40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 254 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  4306   Wed Feb 16 02:04:11 2011 kiwamuUpdateASCIR beam alignment

[Jenne and Kiwamu]

 This time we aligned the vertical angle (not the translation) of the IR beam so that the transmitted light from BS shoots the center of ETMY.

The idea is to use ETMY as a beam pointing reference instead using IP_ANG, assuming the translation is not so bad.

As a result it looks like we are wining. A quick A2L test on ITMX_PITCH showed a small off-centering at sub-milimeter level.

 

 We are concluding that the initial beam after PZT2 had been pointing downward somehow.

Before doing this whole job, we checked the spot shape on IP_POS to see if the beam is clipped or not. It was a round shape, which means no clipping around MMT.

But on the other hand, the spot on IP_ANG had been clipped more than half of its bottom as Suresh reported on his elog (see here).

I found that this clipping is able to be fixed by moving the beam angle upward. I guess the clipping happened at one of the steering mirror in the ETMY chamber.

According to these information, we imagined that the beam was somehow pointing downward after PZT2.

So we started aligning the beam by touching only PZT2 for vertical direction. Then we found a beam spot on ETMY's suspension frame, and brought it to the center.

Then we aligned BS and X arm for this new beam axis. The it resulted a small off-centering on pitch.

Once the MC fully gets back, we will examine the TRX degradation with this configuration.

  4305   Wed Feb 16 01:03:59 2011 JenneUpdateIOOMC alignment work

So.... Kiwamu and I were concerned (still a little concerned) that ETMY is not damping as nicely as it should be.  (It's fine, but the UL rms is ~5, rather than ~1 or less. BURT restores by Kiwamu didn't change anything.) Anyhow, I was heading out to push the annoying ribbon cables more firmly into the satellite adapter board things that are tied to the racks in various places (The back of 1X5 for the corner optics and the end station racks for the ETMs).  The point was to push in the ETMY one, but while I was out in the lab and thinking about it, I also gave all of the corner connectors (MC1, MC2, MC3, ITMx, ITMY, BS, PRM, SRM) a firm push. 

Kiwamu noticed that when I did this, the Mode Cleaner alignment got a little bit worse, as if the connection to the satellite adapter boards hadn't been great, I pushed the connectors in and the connection got better, but we also got a bit of a DC offset in the MC alignment.  Anyhow, the MC_TRANS power went down by ~2, to about the place it had been before Kiwamu adjusted the position of the lens in between the zigzag mirrors.  (I don't know if Kiwamu elogged it earlier, but he scooted the lens a teensy bit closer in the optical path to the Mode Cleaner). 

To counteract this loss in MC transmitted power as a result of my connector actions, I went back to the PSL table and fiddled with the zigzag steering mirrors that steer the beam from the PSL table over to the mode cleaner.  I got it a little better, but it's still not perfect.

Kiwamu has noted that to improve the mode matching into the Mode Cleaner with the new PMC in place, we might have to move the lens which is currently between the zigzag steering mirrors, and put it after the second mirror (so in between the last steering mirror and the pickoff window that sends a piece of the beam over to PSL_POS and PSL_ANG).  This will make the waist between MC1 and MC3 tighter. 

Moral of the story:  To improve IMC mode matching we need to move the last lens closer in the optical path to the mode cleaner waist. Twiddle with zigzag steering mirrors to optimize.

  4304   Tue Feb 15 21:45:08 2011 ranaUpdateIOOThe MC TRANS Story

I forgot to elog that last night I touched up the MC2_TRANS QPD setup. I was perplexed by it always going out of alignment so I investigated.

I found that the fork clamp for the steering mirror for the QPD was not tightened. Shame. The beam diameter was equal to the aperture of the QPD and was clipping. Double shame.

I added a lens and tightened the mounts and centered the beam at ~9 PM yesterday. You can see in the attached trend that the measured power went up by ~10%.

Later, there's a big gap where Valera and Steve change out the PMC. You can see that the MC REFL voltage goes from 4.5 V to 5 V (10% increase in the power delivered to the MC).

There's essentially no change in the total transmission - this indicates that although the PMC transmission is now higher by ~10%, the matching to the IMC has been degraded by an equivalent fraction.

Needs some mode matching work.

Attachment 1: a.png
a.png
  4303   Tue Feb 15 17:48:22 2011 JenneUpdateIOOMode Cleaner resonating again

[Valera, Jenne]

After Steve and Valera switched out the PMC, the Mode Cleaner resonance needed to be brought back.  We spent some time playing with the 2 steering mirrors directly after the PMC, to get the beam through the EOM, and to achieve flashing in the MC.  Valera then adjusted those 2 steering mirrors to minimize MC_REFL_DC.  I did a little bit more, and it's kind of close now, but we're only at ~half normal transmitted power.  Since the 2 steering mirrors after the PMC are so close together, the beam alignment is pretty sensitive to even small touches.  So it's probably time to move on to using the last zigzag steering mirrors on the PSL table, since they're farther apart. 

I have to head out for a little while, but I'll be back in a few hours. Kiwamu said he might continue the alignment into the MC, if he needs the IFO.  Also, we should measure the power before and after the EOM, just to confirm that we're getting through it optimally.  The beam looks good after the EOM, and the MC is resonating, so it should be fine, but it can't hurt to check.

  4302   Tue Feb 15 15:06:25 2011 josephbUpdateCDSCDS todo list for tomorrow morning

Currently, there is a test directory called /opt/rtcds/caltech/c1/new_core where we have the latest svn checkout.  Tomorrow (after everything works), it will become the core directory.

1) Modify on the fb machine the /diskless/root/etc/ld.so.cache file.  This is done by logging into fb, going to /etc/ld.so.conf.d/, modifying epics-x86_64.conf to only have .10 stuff , and running sudo /sbin/ldconfig.  Copy the newly generated /etc/ld.so.cache file to /diskless/root/etc/.

2) Modify the rc.local file on the fb machine in /diskless/root/etc/ to take advantage of the new subscripts and init.d/ start scripts.

3) Add the no_rfm_dma to all the iop models (c1x01,c1x02,c1x03,c1x04,c1x05).

4) Rebuild all front end models with new code.  Install.

5) Build awgtpman and mx_streams with new code.

6) Rerun activateDaq.py (to fix channel names from all the rebuilt code).

7) Double check Burt request files have the switch fix.

8) Restart the front ends.

9)Restart the frame builder.

9) Check channels, exitations, RFM connections.

10) Check Monit is working.

  4301   Tue Feb 15 11:57:06 2011 steve, valeraConfigurationPSLPMC swap

 We swapped the PMC s/n 2677 for s/n lho006.

The table below summarizes the power levels before and after the PMC swap.

  old new
Ptrans 1.32 W 1.42 W
Transmission 85 % 91.5 %
Refl PDDC locked/unlocked 5.0 %  4.3 %
Loss 7-8 % 2-3 %
Leakage out of the back 10 mW 0.3 mW

 

- The power into the PMC (1.67 W) was measured with Scietech bolometer before the first steering PMC mirror. The leakage through the steering mirrors was measured with Ophir power meter to be 12+8 mW. There is also a lens between the mirrors which was not measured. 

- The power through the PMC was measured after the doubler pick off (105 mW), steering mirror (4 mW), and lens (not measured).

- The estimated reflection from four lens surfaces is 1-2% hence 1% uncertainty in the losses in the table.

- The beams into the PMC and on REFL PD were realigned. The beams downstream of the PMC are blocked as we did not realigned the PMC and doubler paths.

- The trans PD ND filters were removed. The VDC=1.28 V now.

- The NPRO current is 2.102 A

 

Atm 1 old

Atm2  new

Attachment 1: P1070421.JPG
P1070421.JPG
Attachment 2: P1070423.JPG
P1070423.JPG
  4300   Tue Feb 15 11:56:17 2011 josephbUpdateCDSUpdated some DAQ channel names
That is my fault for not running the activateDAQ.py script after a round of rebuilds. I have run the script this morning, and confirmed that the oplev channels are showing up in dataviewer.

Quote:

Although Joe and Kiwamu claim that they have inserted the correct DAQ names for the OPLEVs (e.g. PERROR and YERROR) back in Jan. 11, when I look today, I see that these channels are missing!

I want my PERROR/YERRORs back!

 

 

  4299   Tue Feb 15 11:53:02 2011 kiwamuSummaryGeneraltoday's missions

I think this week is going to be an "alignment week".

The goal is to get a good alignment on X arm for both the green and the IR beam in order to reduce a2l couplings.

 

Today's missions are :

 - fixing the oplev channel names (see here)

 - fixing the oplev gain issue (see here)

 - engage the oplev servos

 - f2p adjustment

 - make a realtime lockin model for the f2p measurement and the dithering technique

 - alignment of the MC incident beam  (because we installed a new PMC this morning)

 - manual alignment of the IR beam by steering PZT1 and PZT2 (this procedure will be replaced by an automatic way soon)

 - bounce roll filters (see here)

 

  4298   Tue Feb 15 11:43:53 2011 JenneUpdateComputersOccasional error with NDS2

Just in case anyone has encountered this / knows how to fix it....

I'm running NDS2 on Rossa, trying to get a bunch of raw data from S5.  I get 10min of data at a time, and it goes through ~200 iterations successfully, and then throws the following error:

Getting new data
Connecting.... authenticate ... done
daq_recv_id: Wrong length read (0)
Error reading writerID in daq_recv_block
Warning: daq_request_data failed
 
??? Error using ==> NDS2_GetData
Fatal Error getting channel data.

Error in ==> getDARMdataTS at 37
oot = NDS2_GetData({...

Error in ==> SaveRawData_H1_DARM at 40
    oot = getDARMdataTS(t0s(ii), strideDuration, srate);

  4297   Tue Feb 15 06:59:57 2011 Suresh, JenneOmnistructureGeneralX end enclosure left open

[Jenne, Suresh]

 Jenne found the X-end table enclosure had been left open.  She replaced the lid on it.

 

  4296   Tue Feb 15 06:15:07 2011 SureshUpdateASCno signal from IP_ANG_Seg1

[Valery, kiwamu, Jenne, Suresh]

    I first interchanged the two QPD's on the Y end table to see if the problem QPD related.  Exchanging the units did not make any difference.  The problem therefore had to be in the cables or the circuit boards in 1X4

    We traced the signals pertaining to the IP_ANG QPD ( "Initial Pointing Beam") using  Jay's wiring diagram (pages 2 and 5 of 7).  We noted that while the signals were available on all Segments till the Monitors (Lemo) on 1X4-2-2A card, two of the lines did not reach the output of the cross connect 1X4-B8.  We checked card to make sure that the signals were indeed reaching the back plane of the 1X4-2 chassis using a D990612 extension board.  The card was found to be okay.  We therefore suspected that the cable (CAB_1X4_?) going from the card to the cross connect 1X4-B8 was faulty.  Indeed visual inspection showed that the crimping of the connector was poor and weight of the cable had put further strain on the crimping.  

   I changed the 64-pin connector on the 1X3-2-2A side of the cable. 

When I connected everything back together the problems persisted. Namely the lines P1-1A  (Segment 1 high) and P1-2C (Segment 2 Low) were floating They were not reaching points 2T and 3T respectively on the output of the cross connect.

   I therefore replaced 1X4-B8 with a similar unit which I found in one of the shelves along the East (Y) arm. 

I then checked with the StripTool to make sure that all the quadrants are showing similar response to a flashlight on the QPD.   All Segments are working fine now. Currently the IR Initial Pointing beam reaches the QPD but is not centered on it. 

I did not attempt to center it since the beam appeared to be clipped and may anyway require repositioning.

JD: We need to meditate on where this beam could be getting clipped.  Suresh and I checked that it's not on the viewport on the beam's way out of the ETMY chamber by seeing that the beam is far away from the edges of the viewport, and also far away from the edges of the black beamtube between the viewport and the table.  Suresh mentioned that the clipping nature of the IP_ANG beam sometimes goes away.  I don't know if this is the same clipping that Kiwamu might be seeing with the main beam, or if this is separate clipping just with the IP beam, after it's been picked off.  I suspect it's the same as what Kiwamu is seeing....maybe when we move PZT1, we clip on one of the MMT mirrors or PZT2??  If this is true, it's a total pain since we might have to vent if we can't steer around it.

Y-end_beam.JPG

 

  4295   Tue Feb 15 03:10:37 2011 kiwamuUpdateASCIR beam alignment for Xarm : TRX reduction

I tried aligning the IR beam axis for the X arm to have good beam centering on ITMX and ETMX.

As a first attempt, I started translating the beam upward by steering PZT1 and PZT2, since the pitch was quite off from the center on ITMX.

As a result I could decrease the pitch off-centering down to about 0.5 mm on ITMY, but on the other hand TRX decreased a lot (by a factor of 4).

I am worrying if something in the central part of IFO might be clipping the beam.

 


(notes)

When I was touching PZT1 and PZT2, I payed attention on IP_ANG so that I don't lose a beam spot on IP_ANG.

As long as the beam is on the IP_ANG QPD, the angle of the beam should not be so much different.

Each time after I touched the PZTs, I realigned ITMX and ETMX to maximize the transmitted light.

In this way I proceeded the alignment by changing the PZT offsets little by little while keeping the X arm locked always.

At the beginning, all the PZT offsets were zero. And at the end of this work they became:

 C1:LSC-PZT1_Y = 1.880

 C1:LSC-PZT2_Y = -1.699

But during this alignment work TRX gradually decreased eventually down to 0.25, which had been 1 at the beginning (TRX is calibrated by dividing it by its maximum power).

Along with this TRX reduction, I found that the optical gain also decreased by a factor of about 5.

This fact has been confirmed by intentionally increasing the filter gain such that the servo oscillates at the UGF.
 

Quote:

The amounts of the X arm's beam off-centering have been measured by the A2L technique.

     - ETMX

         PIT  = -1.61 mm

         YAW =  -0.918 mm

    - ITMX

         PIT = -3.76 mm

        YAW = -2.24 mm

 

  4294   Tue Feb 15 02:13:16 2011 kiwamuSummaryASCa daytime task : small signals on ETMX OL

Rana and I found that the QPD for the optical lever at X end are showing small signals.

At this moment each of the segments exhibits approximately 200 counts when the oplev beam is centered.

These small numbers may be due to the coating of ETMX, but we are not sure.

Probably we have to increase the gain of the QPD depending on situations.

 

So a set of the tomorrow's daytime task is:

   1. check the trend data of the QPD outputs to see how much signals were there in the past.

   2. check the whitening filters to make sure if it's on or off.

   3. If it's necessary, increase the gain of the QPD to have reasonable readouts.

I am going to ask somebody to do this task.

  4293   Mon Feb 14 23:29:04 2011 ranaSummarySUSETMX Fitlers moved around

All of the SUS used to have only 1 filter module for SIDE. They now have 3 filter modules for SIDE just like the other DOFs.

Today I moved the filters around so that the sensor filters are in SDSEN, the servo filters are in SUSSIDE, and the dewhitening for the coil is in SDCOIL.

I noticed along the way that the bounce/roll mode notches for all of the suspensions are still set for the frequencies of the previous suspensions. Suresh has 'volunteered' to find the new frequencies and make the new bandstop filters by looking up the seminal work on this by Dan Busby / Sam Waldman.

  4292   Mon Feb 14 21:59:35 2011 ranaUpdateCDSUpdated some DAQ channel names

Although Joe and Kiwamu claim that they have inserted the correct DAQ names for the OPLEVs (e.g. PERROR and YERROR) back in Jan. 11, when I look today, I see that these channels are missing!

I want my PERROR/YERRORs back!

 

  4291   Mon Feb 14 18:27:39 2011 josephbUpdateCDSBegan updating to latest CDS svn, reverted to previous state

[Joe, Alex]

This morning I began the process of bringing our copy of the CDS code up to date to the version installed at Livingston. The motivation was to get fixes to various parts, among others such as the oscillator part.   This would mean cleaning up front end model .mdl files without having to pass clk, sin, cos channels for every optic through 3 layers of simulink boxes.

I also began the process of using a similar startup method, which involved creating /etc/init.d/ start and stop scripts for the various processes which get run on the front ends, including awgtpman and mx_streams.  This allows the monitor software called monit to remotely restart those processes or provide a web page with a real time status of those processes.  A cleaner rc.local file utilizing sub-scripts was also adapted.

I did some testing of the new codes on c1iscey.  This testing showed a problem with the timing part of the code, with cycles going very long.  We think it has something to do with the code not accounting for the fact that we do not have IRIG-B timing cards in the IO chassis providing GPS time, which the sites do have.  We rely on the computer clock and ntpd.

At the moment, we've reverted to svn revision 2174 of the CDS code, and I've put the previously working version of the c1scy and c1x05 (running on the c1iscey computer) back. Its from the /opt/rtcds/caltech/c1/target/c1x05/c1x05_11014_163146 directory.  I've put the old rc.local file back in /diskless/root/etc/ directory on the fb machine.  Currently running code on the other front end computers was not touched.

  4290   Mon Feb 14 17:16:47 2011 steveUpdateGeneralracks, breakers labeled

The south arm labels were changed from y to x to reflect reality.

So racks, manual disconnects and breakers now have their actual name. The east arm will be changed over tomorrow.

Please remove incorrect labels if you see any !

I can not find 1X4 breaker so it will be  traced.

  4289   Mon Feb 14 15:59:49 2011 JenneUpdateIOOWFS quantum efficiency as a function of angle

[Larisa and Jenne]

A few weeks ago (on the 28th of January) I had tried to measure the quantum efficiency of one quadrant of the WFS as a function of angle.  However, Rana pointed out that I was a spaz, and had forgotten to put a lens in front of the laser.  Why I forgot when doing the measurement as a function of angle, but I had remembered while doing it at normal incidence for all of the quadrants, who knows?

Anyhow, Larisa measured the quantum efficiency today.  She used WFS2, quadrant 1 (totally oil-free), since that was easier than WFS1.  She also used the Jenne Laser (with a lens), since it's more stable and less crappy than the CrystaLasers.  We put a 50 Ohm terminator on the RF input of the Jenne Laser, since we weren't doing a swept sine measurement.  Again, the Ophir power meter was used to measure the power incident on the diode, and the reflected power, and the difference between them was used as the power absorbed by the diode for the quantum efficiency measurement.  A voltmeter was used to measure the output of the diode, and then converted to current as in the quote below. 

Still on the to-do list:  Replace the WFS2 diode.  See if we have one around, otherwise order one.  Align beams onto WFS so we can turn on the servo.

QE = (h*c)/(lambda*e) * (I/P)

Where I = (Volts from Pin1 to GND)/2 /500ohms
P = Power from laser - power reflected from diode.
h, c, e are the natural constants, and lambda is 1064nm.
Also, I/P = Responsivity


Larissa is going to put her data and plots into the elog shortly....

Quote:

Quantum Efficiency Measurement:

I refer to Jamie's LHO elog for the equation governing quantum efficiency of photodiodes: LHO 2 Sept 2009

The information I gathered for each quadrant of each WFS was: [1] Power of light incident on PD (measured with the Ophir power meter), [2] Power of light reflected off the PD (since this light doesn't get absorbed, it's not part of the QE), and [3] the photo current output by the PD (To get this, I measured the voltage out of the DC path that is meant to go to EPICS, and backed out what the current is, based on the schematic, attached). 

I found a nifty 25 pin Dsub breakout board, that you can put in like a cable extension, and you can use clip doodles to look at any of the pins on the cable.  Since this was a PD activity, and I didn't want to die from the 100V bias, I covered all of the pins I wasn't going to use with electrical tape.  After turning down the 100V Kepco that supplies the WFS bias, I stuck the breakout board in the WFS.  Since I was able to measure the voltage at the output of the DC path, if you look at the schematic, I needed to divide this by 2 (to undo the 2nd op amp's gain of 2), and then convert to current using the 499 Ohm resistor, R66 in the 1st DC path.  

I did all 4 quadrants of WFS1 using a 532nm laser pointer, just to make sure that I had my measurement procedure under control, since silicon PDs are nice and sensitive to green.  I got an average QE of ~65% for green, which is not too far off the spec of 70% that Suresh found.

I then did all 8 WFS quadrants using the 1064nm CrystaLaser #2, and got an average QE of ~62% for 1064 (58% if I exclude 2 of the quadrants....see below).  Statistics, and whatever else is needed can wait for tomorrow.

Problem with 2 quadrants of WFS2?

While doing all of this, I noticed that quadrants 3 and 4 of WFS2 seem to be different than all the rest.  You can see this on the MEDM screens in that all 6 other quadrants, when there is no light, read about -0.2, whereas the 2 funny quadrants read positive values.  This might be okay, because they both respond to light, in some kind of proportion to the amount of light on them.  I ended up getting QE of ~72% for both of these quadrants, which doesn't make a whole lot of sense since the spec for green is 70%, and silicon is supposed to be less good for infrared than green.  Anyhow, we'll have to meditate on this.  We should also see if we have a trend, to check how long they have been funny.

 

  4287   Mon Feb 14 12:37:23 2011 kiwamuUpdateASCno signal from IP_ANG_Seg1

It turns out there are no reasonable signal from the segment 1 on the IP_ANG QPD.

For right now I can still use it as a funny QPD, but I absolutely need somebody to check and fix it in a daytime.

Quote:

IP_ANG is supposed to be acquired at c1auxey (east end), but actually it had been at c1auxex (south end).

So I fixed it by editing the db files (i.e. ETMXaux.db and ETMYaux.db). Now it seems working fine.

 

  4286   Mon Feb 14 11:23:24 2011 steveUpdateSAFETYenclosure door upgrade

The #3 enclosure door was removed for the carpenter shop this morning. They will make 5 full size doors with window on each.

TEMPORARY aluminum sheet is clamped into its place. This sheet should not be removed till new doors arrive.

The inside Formica color will be changed from white to black to reduce the green scattering.

Attachment 1: P1070418.JPG
P1070418.JPG
  4285   Mon Feb 14 07:58:43 2011 SureshUpdateCDSToday's CDS problems

 

  I am concentrating on the RF system just now and will be attending to the RF PDs one by one.  Also plan to work on some of the simpler CDS problems when I overlap with Joe.  Will be available for helping out with the beam alignment.

 

 

  4284   Mon Feb 14 07:37:13 2011 SureshUpdateElectronicsRF Generation Box: capacitors across power lines

 

There were several parts in this box which did not have shunting capacitors across their input power lines.  Only the four RF amps (ZHL-2) had them.

I soldered two capacitors (100 microF electrolytic and 150pF dipped mica) across the power supply lines of each of the following units:  11MHz oscillator, 29.5 MHz oscillator,  Wenzel 5x frequency multiplier and the 12x RF amplifier (ZHL-1HAD).

It was quite difficult to reach the power inputs of these units as some of them were very close to the inner walls of the box.  To access them I undid the front panel and found that there were several very taut RF cables which prevented me from moving the front panel even a little.

I had to undo some of the RF cables and swap them around till I found a solution in which all of them had some slack.  At the end I checked to make sure that the wiring is in accordance with the schematic present here.

 

  4283   Mon Feb 14 01:40:14 2011 KojiOmnistructureCDSName of the green related channels

I propose to use C1:ALS-xxx_xxx for the names of the green related channels, instead of GCV, GCX, GCY, GFD...

Like C1:SUS or C1:LSC, we name the channels by the subsystems first, then probably we can specify the place.

We can keep the names of the processes as they are now.

  4282   Mon Feb 14 01:19:18 2011 ranaUpdateCDSToday's CDS problems

This is just a listing of  CDS problems I still notice today:

  1. MC2-MCL button was left ON due to BURT failure. This, of course, screws up our Green locking investigations because of the unintended feedback. Please fix the BURT/button issue.

  2. The GCV - FB0 status is RED. I guess this means there's something wrong? Its really a bad idea to have a bunch of whited out or falsely red indicators. No one will ever use these or trust these in the future.
  3. MC1/2/3 Lockins are all white. Also, the MODE switches for the dewhitening are all white.
  4. Is the MC SIDE coil dewhitening filter synced with anything? It doesn't seem to switch anything. Maybe the dewhite indicators at the top right of the SUS screens can be made to show the state of the binary output instead of just the digital filter
  5. MC WFS is all still broken. We need a volunteer to take this on - align beams, replace diodes, fix code/screens.
  4281   Mon Feb 14 00:39:21 2011 rana, sureshUpdateElectronicsVCO Frequency Noise Measurement with the MFD

We hooked up the VCO Driver output to the MFD. We adjusted the levels with attenuators to match up to the Level 7 mixer that's being used.

The mixer the input to the SR560 is going in to the XARM_COARSE_OUT channel and the SR560 (AC coupled, Low Noise, G=1000, LP@1kHz) 600 Ohm output goes into XARM_FINE_OUT.

We calibrated these channels by putting in a 10 mVpp sine wave at 0.22 Hz into the Wideband Input of the VCO Driver box (which has been calibrated to have 1.75 MHz/V for f < 1.6 Hz). This should correspond to 17.5 kHz_pp.

To increase the sensitivity, we also added a 140 ft. BNC cable to the setup. We also added some extra short cable to make the overall phase shift be ~90 deg and zero out the mixer output.

I used the time series data in DTT to then calibrate the channels by changing the GAIN field in their filter modules. So now the DAQ channels are both calibrated as 1 count/Hz.

 

  4280   Sun Feb 13 16:50:17 2011 kiwamuUpdateASCIP_ANG was at wrong ADC

I found that the ADC channels for IP_ANG had been assigned to a wrong machine.

IP_ANG is supposed to be acquired at c1auxey (east end), but actually it had been at c1auxex (south end).

This is the reason why we couldn't see any signals from IP_ANG.

So I fixed it by editing the db files (i.e. ETMXaux.db and ETMYaux.db). Now it seems working fine.

 

This mistake obviously came from the X-Y name swapping business. Something else might be still wrong.

xmen.jpg

 

  4279   Sun Feb 13 16:20:53 2011 SureshUpdateElectronicsREFL11 Photodiode replaced

It is 0.375 mW as in the calculation.  The total diode output is just 1mW and it is divided with a 50/50 beam splitter...  There are a couple of lenses along the way which may account for the ~12% loss. 

I used a handheld multimeter to measure the output.  

 

  4278   Sun Feb 13 15:02:23 2011 kiwamuUpdateGreen LockingX arm beam offcentering has been measured

The amounts of the X arm's beam off-centering have been measured by the A2L technique.

So now we are able to start aligning the IR beam axis in a quantitative way.

 

(motivation)

 Since we saw big residual motions at 1 Hz, 16 Hz on both the green beat note signal and the IR PDH signal (see #4268 and #4211),

we are suspecting that these noise come from an angle to length coupling.

In order to minimize the angle to length coupling, one thing we can do is to bring the beam spots to the center of ITMX and ETMX more precisely.

To do it, we have to quantitatively know how well the beam spots are on the center of the optics. Therefore I started measuring the amount of the beam off-centering.

 

(method)

 The A2L technique was used to measure the off-centering with the real-time lockin system, which has been recently embedded in the real-time code by Joe (see #4265).

The idea is the same as Yuta did before (see #3863).

But this time the excitation signal from the real-time oscillator was injected directly to the coil matrix on either ITMX or ETMX, at 18.13 Hz with the amplitude of about 400 cnt.

When the IR laser stays locked to the X arm, the LSC feedback signal is demodulated with the oscillator signal.

This demodulated signal gives the amount of the off-centering.

For this purpose I modified Yuta's A2L script such that we can use it also for the X arm.

 

(results)

 I obtained the following values:

     - ETMX

         PIT  = -1.61 mm

         YAW =  -0.918 mm

    - ITMX

         PIT = -3.76 mm

        YAW = -2.24 mm

I used the same calibration factor as that of Koji calculated (see #3020) for MC, in order to convert the results from the coil gain to the off-centering.

These values are consistent with the spots appearing on the CCD monitors.

 misposition.png

  4277   Sun Feb 13 02:33:37 2011 KojiUpdateElectronicsREFL11 Photodiode replaced

Suresh is saying 375mW and 0.375mW. Let's wait for his update of the actual power.

Also he is not using EPICS, there may be the factor of two missing for now.

Quote:

I also checked to see if we have a DC output from the new PD.  With 375mW of 1064nm light incident we have 15mV of output.  Which matches well with the typical Reponsivity of 0.8V/A reported in the datasheet and our REFL11 ckt .  The schematic of the ckt is also attached here for easy reference.  The various factors are

V_dc = 0.375 mW x 0.8 V/A x 10 Ohm x 5 = 15mV

 

  4276   Sat Feb 12 23:22:21 2011 ranaUpdateElectronicsREFL11 Photodiode replaced

375 mW is way too much light. We must never put more than 100 mW on any of these diodes. We don't want to blow up more diodes like we did with the WFS. The InGaAs diodes often show an excess dark noise before they finally let go and completely fail. This one may show excess during the shot noise testing.

We should ensure that the beam paths are engineered so that none of these new detectors ever sees such high light levels.

The DC path should be made to let us see a 10V from the differential EPICS readout when there is 100 mA of photocurrent (i.e. an effective 100 Ohms transimpedance):

0.1 A  * 10 V/A * 5 V/V * 2V/V

The last factor of 2 is from the single to differential conversion.

If we really only get 15 mV from 375 mW, then this diode or the circuit is broken.

  4275   Sat Feb 12 08:08:05 2011 SureshUpdateElectronicsREFL11 Photodiode replaced

A new photodiode ( Perkin and Elmer Model no. C30642GH Sl No.1526) has been installed in the place of the old photodiode.  The datasheet of this model is attached. 

The 68pF capacitor which was present in parallel with the photodiode has been removed.  Here is a picture of the PCB ( in all its gory detail!) and the photodiode after replacement.

  P2120565.JPG   P2120571.JPG

 

I also checked to see if we have a DC output from the new PD.  With 375mW of 1064nm light incident we have 15mV of output.  Which matches well with the typical Reponsivity of 0.8V/A reported in the datasheet and our REFL11 ckt .  The schematic of the ckt is also attached here for easy reference.  The various factors are

V_dc = 0.375 mW x 0.8 V/A x 10 Ohm x 5 = 15mV

The last factor is the gain of the last stage on the DC route.

When I reassembled the box I noticed that there is problem with the SMA connectors popping out of the box.  The holes seem misplaced so I enlarged the holes to remove this concern.

P2120572.JPG

 

Attachment 1: C30642_datasheet-1.pdf
C30642_datasheet-1.pdf C30642_datasheet-1.pdf C30642_datasheet-1.pdf C30642_datasheet-1.pdf C30642_datasheet-1.pdf
Attachment 2: 40mUpgradeREFL11schematic.pdf
40mUpgradeREFL11schematic.pdf
  4274   Fri Feb 11 16:43:09 2011 steveUpdateVIDEOMC1 & 3 video monitor

I set up video monitoring of MC1 and MC3

Attachment 1: P1070415.JPG
P1070415.JPG
  4273   Fri Feb 11 09:27:03 2011 steveConfigurationVACRGA scan

The RGA scan is normal at day 52 of this pump down.

Light power BS 1064nm ~25mW, ETMX 532nm ~5mW

Attachment 1: rgascan20110211.jpg
rgascan20110211.jpg
Attachment 2: pressurepd70.jpg
pressurepd70.jpg
  4272   Fri Feb 11 00:20:58 2011 SureshUpdateElectronicsREFL11: Photodiode requires replacement

 

This is with reference to Kevin and Jenne's elogs  # 3890, 4034 and 4048

While the electronics are working okay, there is no DC signal from the photodiode. 

Since the solderings and tracks on the PCB were fine I took a close look at the exposed front face of the photodiode.

REFL11_10Feb2011.jpg

As we can see, one of the thin wires on the top surface of the photodiode is broken.  We can see some wipe marks closer to the lower left edge..

Something seems to have brushed across the exposed face of the photodiode and dislodged the wire.

 

Question:

The new photodiode still has its protective can intact.   Do we need to remove the can and expose the photodiode before istallation?

 

  4271   Thu Feb 10 14:52:36 2011 AidanHowToComputersAdding filenames in MATLAB plots

The following code is incredibly useful when creating  MATLAB plots as it adds the filename of the script to the plot itself. I think it should be used for all MATLAB plots that go on the elog.

For example, I have no idea where the data/script is that was used to generate these plots.

 

orient landscape

xposn = 0.0;

yposn = -0.13; % you sometimes have to tweak this value depending on the page size and the number of subplots

text(xposn,yposn,[filename('fullpath'), '.m'], ...

     'Units', 'normalized', ...

     'Interpreter', 'none', ...

     'FontSize', 6)

print('-dpdf', [filename('fullpath'), '.pdf'])

  4270   Thu Feb 10 14:07:18 2011 josephbUpdateCDSUpdating dolphin drivers to eliminate timeouts when one dolphin card is shutdown

[Joe,Alex]

Alex came over and we installed the new Dolphin drivers so that the front ends using the Dolphin PCIe RFM network don't pause for a long time when one of the other nodes in the network go down.  Generally this pause would cause the code to time out and quit.  Now you can take c1lsc or c1sus down without having the other have problems.

We did note on reboot however, that the Dolphin_wait script sometimes (not always) seems to hang.  Since this is run at boot up, to ensure the dolphin card has had enough to allocate memory space for data to be written/read from by the IOP process, it means nothing else in the startup script gets run if it does happen.  In this case, running "pkill dolphin_wait" may be necessary.

Note that you may still have problems if you hit the power button to force a shutdown (i.e. holding it for 4 seconds for immediate power off), but as long as you do a "reboot" or "shutdown -r now" type command, it should come down gracefully. 

 

What was done:

Alex grabbed the code from his server, and put it /home/controls/DIS/ on fb.

He ran the following commands in that directory to build the code.

./configure  '--with-adapter=DX' '--prefix=/opt/DIS'

make

sudo make install

He proceeded to modify the /diskless/root/etc/rc.local to have the line:

insmod /lib/modules/2.6.34.1/kernel/drivers/dis/dis_kosf.ko

In that same file he commented out

cd /root

and

exec /bin/bash/

He then modified the run levels in /diskless/root/etc/inittab. Level 0, level 3, and level 6 were changed:

l0:0:wait/etc/rc.halt

l3:3:wait:etc/rc.level3

l6:6:wait:/etc/rc.reboot

Then he created the scripts he was refering to:

rc.level3 is just:

exec /bin/bash

rc.halt is:

/opt/DIS/sbin/dxtool prepare-shutdown 0
sleep 3
halt -p

rc.reboot is:

reboot

 

Basically rc.halt calls a special code which prepares the Dolphin RFM card to shutdown nicely.  This is why just hitting the power button for 4 seconds will cause problems for the rest of the dolphin network.

We then checked out of svn the latest dolphin.c in  /opt/rtcds/caltech/c1/core/advLigoRTS/src/fe

 

The Dolphin RFM cards have a new numbering scheme.  4 is reserved for special  broadcasts to everyone, so the Dolphin node IDs now start at 8.  So we needed to change the c1lsc and c1sus Dolphin node IDs.

To change them we went to /etc/dis/dishosts.conf on the fb machine, and changed the following lines:

 

HOSTNAME: c1sus
ADAPTER:  c1sus_a0 4 0 4
HOSTNAME: c1lsc
ADAPTER:  c1lsc_a0 8 0 4

to

HOSTNAME: c1sus
ADAPTER:  c1sus_a0 8 0 4
HOSTNAME: c1lsc
ADAPTER:  c1lsc_a0 12 0 4

The FE models for the c1lsc and c1sus machines were recompiled and then the computers were rebooted.  After having them come back up, we tested that there was no time out by shutting down c1lsc and watching c1sus. We then reveresed and shutdown c1sus while watching c1lsc.  No problems occured.  Currently they are up and communicating fine.

 

  4269   Thu Feb 10 11:16:31 2011 steveUpdatePEMsouth arm AC turned on

The air condition was off for the south arm. I  turned it on.

  4268   Thu Feb 10 05:06:35 2011 kiwamuUpdateGreen Lockingbeat noise : a little bit better, and 1Hz peak from amplitude noise coupling

 I repeated the same measurement as that Koji did before (see here) with the mixer-based frequency discriminator.

The frequency fluctuation of the beat note is now 50 kHz in rms integrated down to 0.1 Hz, which is a bit better than before.

However there still is the same undesired structure in the spectrum below 10 Hz.

 

 Screen_shot_2011-02-10_at_4.01.43.png

Fig.1 power spectra of the green beat note fluctuation in terms of frequency fluctuation.

   Red curves were taken when the IR was locked to the MC, and the green was locked to the X arm.

Blue curves were taken when both the IR and the green were locked to the X arm.

Black curve was also the one taken when the IR and the green were locked to the X arm, but showing the lower noise level.

I have no idea what exactly was going on when I took the black curve, but this noise level sometimes showed up.

The discrepancy may come from a kind of calibration error although I kept using the same calibration factor to convert the data from count to frequency.

Need more investigations.

 


 Additionally Koji and I took the coherence between the beat fluctuation and the transmitted lights of both the IR and the green.

It showed a strong coherence at 1 Hz, which is one of the dominant noise of the beat note.

This probably indicates that the 1 Hz peak is produced by a coupling from amplitude fluctuation.

Screen_shot_2011-02-10_at_4.52.13.png

 For monitoring the green transmitted light, I used the Jenne's PD (see here)

  4267   Thu Feb 10 00:23:25 2011 JenneUpdateGreen LockingGreen TRX DC PD installed on PSL

Using a stray beam that is generated as the transmitted green beam from the Xarm goes through the viewport to the PSL table, I installed a fast lens (because I was constrained for space) and a Thorlabs PDA36 photodiode on the PSL table.

The BNC cable runs along the edge of the PSL table, up the corner hole with the huge bundle of cables, and over to IOO_ADC_0. It's channel 3 on the simulink model, which means that it is plugged into connector #4.

With the green resonating TEM00, I have ~1.4V output from the photodiode, as seen on a voltmeter. This corresponds to ~1500 counts on the MEDM screen.

 

Note to self:  Switch to a ~1cm diode with a boatload of gain (either from the 40m or Bridge), and use transmission through a steering mirror of the actual beat note path, not the jittery viewport pickoff.  Want RIN noise level to be about 1e-5, only care about below ~100Hz so don't need broadband.

  4266   Wed Feb 9 23:48:12 2011 SureshConfigurationCamerasVideo Cable work: New Labels

[Larisa, Aidan,Steve,Suresh]

   Today was the first session for implementing the new video cabling plan laid out in the document " CCD_Cable_Upgrade_Plan_Jan11_2011.pdf"  by Joon Ho attached to his elog entry 4139.  We started to check and label all the existing cables according to the new naming scheme. 

So far we have labeled the following cables. Each has been checked by connecting it to a monitor near the Video Mux and a camera at the other end.

C1:IO VIDEO 8ETMYF

C1:IO-VIDEO 6 ITMYF

C1:IO-VIDEO 21 SRMF

C1:IO-VIDEO 25 OMCT

C1:IO-VIDEO 19 REFL

C1:IO-VIDEO 22 AS

C1:IO-VIDEO 18 IMCR

C1:IO-VIDEO 14 PMCT

C1:IO-VIDEO 12 RCT

C1:IO-VIDEO 9 ETMXF

C1:IO-VIDEO 1 MC2T

 

Next we need to continue and finish the labeling of existing cables.  We then choose a specific set of cables which need to be laid together and proceed to lay them after attaching suitable lables to them.

 

 

  4265   Wed Feb 9 15:26:22 2011 josephbUpdateCDSUpdated c1scx with lockin, c1gcv for green transmission pd

Updated the c1scx model to have two Lockin demodulators (C1:SUS-ETMX_LOCKIN1 and C1:SUS-ETMX_LOCKIN2).  There is a matrix C1:SUS-ETMX_INMUX which directs signals to the inputs of LOCKIN1 and LOCKIN2.  Currently only the GREEN_TRX signal is the only signal going in to this matrix, the other 3 are grounds.  The actual clocks themselves had to be at the top level (they don't work inside blocks) and thus named C1:SCX-ETMX_LOCKIN1_OSC and C1:SCX-ETMX_LOCKIN2_OSC.

 

There is a signal (IPC name is C1:GCV-SCX_GREEN_TRX) going from the c1gcv model to the c1scx model, which will contain the output from Jenne's green transmission PD which will eventually be placed. I've placed a filter bank on it in the c1gcv model as a monitor point, and it corresponds to C1:GCV-GREEN_TRX.

 

The suspension control screens were modified to have a screen for the Matrix feeding signals into the two lockin demodulators.  The green medm screen was also modified to have readbacks for the GREEN_TRX and GREEN_TRY channels.

 

So on the board, the top channel (labeled 1, corresponds to code ADC_0_0) is MCL.

Channel 2 (ADC_0_1) is assigned to frequency divided green signal.

Channel 3 (ADC_0_2) is assigned to the beat PD's DC output.

Channel 4 (ADC_0_3) is assigned to the green power transmission for the x-arm.

Channel 5 (ADC_0_4) is assigned to the green power transmission for the y-arm.

  4264   Wed Feb 9 10:25:46 2011 steveUpdateSAFETYAP table

I blocked the  AP table's south west 10" ID port since it is obsolete with the new layout.

Reminder: items on the enclosure self can fall down in an earthquake. I moved oscilloscope and heavy calorimeter head from the edge of the cliff.

Attachment 1: P1070395.JPG
P1070395.JPG
  4263   Tue Feb 8 16:44:43 2011 JenneUpdateComputersLIGO Grid Cluster client upgraded on Rossa

I did a yum-install of the latest ldg-client (to get onto the LIGO Clusters) on Rossa. 

I followed the instructions on the wiki page, and everything seemed to work nicely.

I think the new ldg client installs somewhere on the local computer, so if anyone wants cluster access on any other computer, they should follow the same directions.

  4262   Tue Feb 8 16:04:58 2011 josephbUpdateCDSHard coded decimation filters need to be fixed

[Joe, Rana]

Filter definitions for the decimation filters to epics readback channels (like _OUT16) can be found in the fm10Gen.c code (in /opt/rtcds/caltech/c1/core/advLigoRTS/src/include/drv).

At the moment, the code is broken for systems running at 32k, 64k as they look to be defaulting to the 16k filter.  I'd like to also figure out the notation and plot the actual filter used for the 16k.

Rana has suggested a 2nd order, 2db ripple low pass Cheby1 filter at 1 Hz.

 

  51 #if defined(SERVO16K) || defined(SERVOMIXED) || defined(SERVO32K) || defined(SERVO64K) || defined(SERVO128K) || defined(SERVO256K)
  52 static double sixteenKAvgCoeff[9] = {1.9084759e-12,
  53                                      -1.99708675982420, 0.99709029700517, 2.00000005830747, 1.00000000739582,
  54                                      -1.99878510620232, 0.99879373895648, 1.99999994169253, 0.99999999260419};
  55 #endif
  56
  57 #if defined(SERVO2K) || defined(SERVOMIXED) || defined(SERVO4K)
  58 static double twoKAvgCoeff[9] = {7.705446e-9,
  59                                  -1.97673337437048, 0.97695747524900,  2.00000006227141,  1.00000000659235,
  60                                  -1.98984125831661,  0.99039139954634,  1.99999993772859,  0.99999999340765};
  61 #endif
  62
  63 #ifdef SERVO16K
  64 #define avgCoeff sixteenKAvgCoeff
  65 #elif defined(SERVO32K) || defined(SERVO64K) || defined(SERVO128K) || defined(SERVO256K)
  66 #define avgCoeff sixteenKAvgCoeff
  67 #elif defined(SERVO2K)
  68 #define avgCoeff twoKAvgCoeff
  69 #elif defined(SERVO4K)
  70 #define avgCoeff twoKAvgCoeff
  71 #elif defined(SERVOMIXED)
  72 #define filterModule(a,b,c,d) filterModuleRate(a,b,c,d,16384)
  73 #elif defined(SERVO5HZ)
  74 #else
  75 #error need to define 2k or 16k or mixed
  76 #endif

  4261   Tue Feb 8 15:22:13 2011 kiwamuUpdateGreen Lockingnew electronics rack at X end

 Yesterday I moved the whole green electronics stuff, which had been sitting on the floor at the X end,  into a new electronics rack.

The rack now is placed under the cable rail close to the ETMX chamber.

DSC_2861_ss.png

  4260   Tue Feb 8 13:26:11 2011 AidanSummaryGreen LockingTemperature dependence of phase change of green on reflection

 I did a quick back of the envelope calculation of the expected green phase change on reflection from the aLIGO ITM.

The phase change per nm, K1 = delta phi/delta Lambda, around 532nm is ~1.5 degrees/nm (from the LMA data) [this number is approximately 100x smaller at 1064nm]

I assumed that very small changes in the thickness of the coating appear equivalent to shifting the spectra for reflection/transmission/phase-change-on-reflection up or down by delta lambda, where

delta Lambda/Lambda = delta h/h

where h is the total thickness of the coating and delta h is the change in the thickness of the coating.

Assume that delta h/h = alpha deltaT, where alpha is the coefficient of thermal expansion and delta T is the change in temperature. (approximately 1K)

Then delta phi = K1* Lambda * alpha * delta T = 1.5 degrees/nm * 532nm * 10^-5 K^-1 * 1.0 K =  8 * 10^-3 degrees.

Assume that 360 degree phase change corresponds to one FSR.

Therefore, the frequency shift due to temperature change in the coating = 8*10^-3/360 * FSR = 2.2 *10^-5 * FSR.

Therefore, the expected frequency shift per degree temperature change = 2.2*10^-5 * FSR [Hz/K]

  4259   Tue Feb 8 10:23:02 2011 AidanSummaryGreen LockingDigital Frequency Discriminator - reference

 

Here's the reference for the self-reference frequency detection idea. See Figure 2.

http://www.phys.hawaii.edu/~anita/new/papers/militaryHandbook/mixers.pdf

  4258   Mon Feb 7 21:23:11 2011 ranaConfigurationPSLPSL FSS Temperature Sensor Interface box removed

I noticed that the RMTEMP channel was spiking myteriously when Kiwamu opened the PSL door. We found out that the LEMO connectors would intermittently short to the case and cause ~1 deg steps in the temeprature.

We have removed the case and examined it. Not only were the connections to the box intermittent, there was a cold solder joint inside on an unsecured flying add-on opamp. The whole thing is a giant hack.

PK was the last person to work on this box, but I'm sure that he wouldn't have left it in this state. Must be gremlins.

P2070555-1.JPG

The LEMO connectors on the front are the ones touching. The LT1021 is the badly soldered part.

  4257   Mon Feb 7 19:21:32 2011 Beard PapaMetaphysicsPhotosThe Adventures of Dr Stochino and Beard Papa

  4256   Mon Feb 7 10:37:28 2011 josephbUpdateComputersTemporarily removed cronjob for rsync.backup

The backup appears to have finished on nodus, and I've put the rsync job back in the crontab.

Quote:

I removed the rsync backup from nodus' crontab temporarily so as to not have multiple backup jobs running.  The job I started from yesterday was still running.  Hopefully the backup will finish by Monday.

The line I removed was:

0 5 * * * /opt/rtcds/caltech/c1/scripts/backup/rsync.backup

ELOG V3.1.3-